# Phenomenology of decaying turbulence beneath surface waves

### v2

@@ Abstract
This paper explores decaying turbulence beneath surface waves that is initially isotropic and shear-free. We start by presenting phenomenology revealed by wave-averaged numerical simulations: an accumulation of angular momentum in coherent vortices perpendicular to the direction of wave propagation, suppression of kinetic energy dissipation, and the development of depth-alternating jets. We interpret these features through an analogy with rotating turbulence (Holm, 1996), wherein the curl of the Stokes drift, ×$\bm{\nabla}\times\bm{u}^{\mathrm{S}}$, takes on the role of the background vorticity (for example, $(f_{0}+\beta y)\hskip 1.00006pt\bm{\hat{z}}$ on the beta plane). We pursue this thread further by showing that a two-equation model proposed by Bardina et al. (1985) for rotating turbulence reproduces the simulated evolution of volume-integrated kinetic energy. This success of the two-equation model — which explicitly parametrizes wave-driven suppression of kinetic energy dissipation — carries implications for modeling turbulent mixing in the ocean surface boundary layer. We conclude with a discussion about a wave-averaged analogue of the Rossby number appearing in the two-equation model, which we term the “pseudovorticity number” after the pseudovorticity ×$\bm{\nabla}\times\bm{u}^{S}$. The pseudovorticity number is related to the Langmuir number in an integral sense.
@@ 2 Simulations of decaying turbulence beneath surface waves
where $\bm{u}^{\mathrm{S}}=u^{\mathrm{S}}(z)\hskip 1.00006pt\bm{\hat{x}}$ is the surface wave Stokes drift (Craik & Leibovich, 1976; Suzuki & Fox-Kemper, 2016; van den Bremer & Breivik, 2018; Vanneste & Young, 2022). We refer to ×$\bm{\nabla}\times\bm{u}^{\mathrm{S}}$ — the curl of the surface-wave pseudomomentum $\bm{u}^{\mathrm{S}}$ (Andrews & McIntyre, 1978) — as the “pseudovorticity”. In the context of equation (1), the main difference between rotating turbulence and turbulence beneath steady surface waves is the spatial structure of $\partial_{z}u^{\mathrm{S}}$: in shallow water $\partial_{z}u^{\mathrm{S}}\sim z$ to leading-order, while $\partial_{z}u^{\mathrm{S}}$ decays exponentially for deep water waves.
@@ 2.1 Coherent structures in rotating and wave-averaged turbulence
[…]
@@
Figure 2: As figure 1 but showing (a) $\zeta$ for rotating turbulence in the $xy$-plane; (b) $\eta$ for turbulence beneath surface waves in the $xz$-plane, (c) $\eta$ for isotropic turbulence in the $xz$-plane.
@@ 2.2 Zonation and the analogy of wave-averaged turbulence with beta-plane turbulence
Figure 3: The evolution of cross-wave momentum $v$ in the $xz$-plane (6 panels on the left) and horizontally-averaged along-wave-momentum $u$ (2 panels on the right) at $t=40,400$ and for three wave fields: “deep” (left panel, red lines), “medium” (middle panel, orange lines), and “weak” (right panel, blue lines). Dashed lines in the top right plot show the Stokes drift profile (normalized) for each case. The “medium” waves case uses the background vorticity $\bm{\Omega}_{\text{waves}}$ in (4). Both the $u$ profiles and the light gray “zero lines” are spaced apart by the scale $\delta u=10^{-2}$. The $u$-profiles at right show the development of depth-alternating jets, including a counter-wave surface jet for all cases. The $v$-slices are similar between the three cases at early times, but at later times exhibit strong, localized wave-impacts in their respective regions of significant Stokes shear. Note that $v$ plays the role that vertical velocity plays in rotating turbulence. Note that the Rhines scale may be estimated as $\ell_{R}=2\pi\sqrt{U/\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}}$. With $\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}=\tfrac{1}{2}$ and $U\approx 10^{-2}$ for the medium waves case, for example, we estimate $\ell_{R}\approx 0.8$, consistent with the jet spacing apparent on the right panels.
@@ 3 Phenomenological model for the evolution of kinetic energy
where $n=1/(a-1)=6/5$. (See also equation (10) in Bardina et al. (1985).) Taking the limit $\Omega\to 0$ (or $t\to 0$ with finite $\Omega$), we obtain the corresponding solution for isotropic turbulence,
@@ 4.1 The “pseudovorticity” number
The analogy with rotating turbulence inspires the definition of a new non-dimensional number for characterizing surface wave effects on turbulence. In rotating turbulence, the Rossby number measures the relative magnitude of the relative vorticity ×$\bm{\nabla}\times\bm{u}$ and the background vorticity, $f\hskip 1.00006pt\bm{\hat{z}}$, such that
@@ 4.2 Connection with the Langmuir number
bulk $\text{bulk Ps}=\frac{W}{\Omega H}\sim\frac{u_{\star}}{u^{\mathrm{S}}(z=0)}\sim\mathrm{La^{2}}\,.$ (23)
@@
[Author ORCIDs]Gregory L. Wagner, $0$-0001-5317-2445; Navid C. Constantinou, $0$-0002-8149-4094.

### v3

@@ Abstract
This paper explores decaying turbulence beneath surface waves that is initially isotropic and shear-free. We start by presenting phenomenology revealed by wave-averaged numerical simulations: an accumulation of angular momentum in coherent vortices perpendicular to the direction of wave propagation, suppression of kinetic energy dissipation, and the development of depth-alternating jets. We interpret these features through an analogy with rotating turbulence (Holm, 1996), wherein the curl of the Stokes drift, $\bm{\nabla}\times\bm{u}^{\mathrm{S}}$, takes on the role of the background vorticity (for example, $(f_{0}+\beta y)\hskip 1.00006pt\bm{\hat{z}}$ on the beta plane). We pursue this thread further by showing that a two-equation model proposed by Bardina et al. (1985) for rotating turbulence reproduces the simulated evolution of volume-integrated kinetic energy. This success of the two-equation model — which explicitly parametrizes wave-driven suppression of kinetic energy dissipation — carries implications for modeling turbulent mixing in the ocean surface boundary layer. We conclude with a discussion about a wave-averaged analogue of the Rossby number appearing in the two-equation model, which we term the “pseudovorticity number” after the pseudovorticity $\bm{\nabla}\times\bm{u}^{S}$. The pseudovorticity number is related to the Langmuir number in an integral sense.
@@ 2 Simulations of decaying turbulence beneath surface waves
where $\bm{u}^{\mathrm{S}}=u^{\mathrm{S}}(z)\hskip 1.00006pt\bm{\hat{x}}$ is the surface wave Stokes drift (Craik & Leibovich, 1976; Suzuki & Fox-Kemper, 2016; van den Bremer & Breivik, 2018; Vanneste & Young, 2022). We refer to $\bm{\nabla}\times\bm{u}^{\mathrm{S}}$ — the curl of the surface-wave pseudomomentum $\bm{u}^{\mathrm{S}}$ (Andrews & McIntyre, 1978) — as the “pseudovorticity”. In the context of equation (1), the main difference between rotating turbulence and turbulence beneath steady surface waves is the spatial structure of $\partial_{z}u^{\mathrm{S}}$: in shallow water $\partial_{z}u^{\mathrm{S}}\sim z$ to leading-order, while $\partial_{z}u^{\mathrm{S}}$ decays exponentially for deep water waves.
@@ 2.1 Coherent structures in rotating and wave-averaged turbulence
Above, $\tfrac{1}{2}z$ is the pseudovorticity associated with a shallow water wave in $z\in[0,1]$ with Stokes drift $\bm{u}^{\mathrm{S}}\approx\tfrac{1}{2}\big{(}1+\tfrac{1}{2}z^{2}\big{)}\hskip 1.00006pt\bm{\hat{x}}$. The domain is horizontally-periodic in $x,y$, and we use free-slip boundary conditions at the top and bottom boundaries in $z$. Implicit dissipation of grid-scale kinetic energy is provided by an upwind-biased, nominally 9th-order Weighted, Essentially Non-Oscillatory advection scheme (Pressel et al., 2017; Shu, 2020; Silvestri et al., 2024). Our implicit dissipation method allows us to simulate decaying turbulence that undergoes a transition from isotropic to nearly two-dimensional dynamics.
@@
Figure 2: As figure 1 but showing (a) $\zeta$ for rotating turbulence in the xyxy-plane; (b) $\eta$ for turbulence beneath surface waves in the $xz$-plane, (c) $\eta$ for isotropic turbulence in the xzxz-plane.
@@ 2.2 Zonation and the analogy of wave-averaged turbulence with beta-plane turbulence
Figure 3: The evolution of cross-wave momentum $v$ in the xzxz-plane (6 panels on the left) and horizontally-averaged along-wave-momentum $u$ (2 panels on the right) at $t=40,400$ and for three wave fields: “deep” (left panel, red lines), “medium” (middle panel, orange lines), and “weak” (right panel, blue lines). Dashed lines in the top right plot show the Stokes drift profile (normalized) for each case. The “medium” waves case uses the background vorticity $\bm{\Omega}_{\text{waves}}$ in (4). Both the $u$ profiles and the light gray “zero lines” are spaced apart by the scale $\delta u=10^{-2}$. The $u$-profiles at right show the development of depth-alternating jets, including a counter-wave surface jet for all cases. The $v$-slices are similar between the three cases at early times, but at later times exhibit strong, localized wave-impacts in their respective regions of significant Stokes shear. Note that $v$ plays the role that vertical velocity plays in rotating turbulence. Note that the Rhines scale may be estimated as $\ell_{R}=2\pi\sqrt{U/\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}}$. With $\big{|}\partial_{z}^{2}u^{\mathrm{S}}\big{|}=\tfrac{1}{2}$ and $U\approx 10^{-2}$ for the medium waves case, for example, we estimate $\ell_{R}\approx 0.8$, consistent with the jet spacing apparent on the right panels.
@@ 3 Phenomenological model for the evolution of kinetic energy
where n=1/(a1)=6/5n=1/(a-1)=6/5. (See also equation (10) in Bardina et al. (1985).) Taking the limit $\Omega\to 0$ (or $t\to 0$ with finite $\Omega$), we obtain the corresponding solution for isotropic turbulence,
@@ 4.1 The “pseudovorticity” number
The analogy with rotating turbulence inspires the definition of a new non-dimensional number for characterizing surface wave effects on turbulence. In rotating turbulence, the Rossby number measures the relative magnitude of the relative vorticity $\bm{\nabla}\times\bm{u}$ and the background vorticity, $f\hskip 1.00006pt\bm{\hat{z}}$, such that
@@ 4.2 Connection with the Langmuir number
$\text{bulk Ps}=\frac{W}{\Omega H}\sim\frac{u_{\star}}{u^{\mathrm{S}}(z=0)}\sim\mathrm{La^{2}}\,.$ (23)
@@
[Author ORCIDs]Gregory L. Wagner, 0000-0001-5317-2445; Navid C. Constantinou, 0000-0002-8149-4094.